OpenStudy (anonymous):

A rectangular sold has a volume of 640 cubic units. The width is 3 units more than the height and the length is 1 unit more than three times the height. Find the dimension

OpenStudy (zpupster):



OpenStudy (zpupster):

can you see it now??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what to do?

OpenStudy (zpupster):

V= (L) (W)(H) plug in oyur values and mult

OpenStudy (anonymous):

V= (1+3H)(H+3)(H)

OpenStudy (anonymous):

@zpupster what to do now?

OpenStudy (zpupster):

mult LHS left ahnd side out

OpenStudy (anonymous):

huh? pls explaim

OpenStudy (zpupster):

do you know the foil method

OpenStudy (anonymous):

no. pls help

OpenStudy (zpupster):

if you have a binomial like this: (a+b)(c+d) = you mult first, outside inside and last terms FOIL

OpenStudy (anonymous):

3h^3+10h^2+3h

OpenStudy (zpupster):

i will get you started: (1+3H)(H+3) First: 1*H Outside:1*3 Inside:3H^2 Last:=?????

OpenStudy (zpupster):

yes good

OpenStudy (anonymous):

then? what to do?

OpenStudy (zpupster):

subtract 640 from both sides

OpenStudy (anonymous):

how?

OpenStudy (zpupster):

3h^3+10h^2+3h -640 = 640 - 640

OpenStudy (anonymous):

and then?

OpenStudy (zpupster):

factor the LHS

OpenStudy (zpupster):

(H-5)(128 +25+3h^2) = 0

OpenStudy (zpupster):

split into 2 equations and solve for H (H-5)=0 (128 +25+3h^2)=0

OpenStudy (anonymous):

i think its (128+24+3h^2)=0?

OpenStudy (anonymous):

@hartnn pls help

OpenStudy (zpupster):

ok solve for the first equation first

OpenStudy (anonymous):

H=5

OpenStudy (zpupster):

128+25+3h^2= 0 divide by 3 128/3 + 25/3 +h^2 = 0 then subtract 128/3 from both sides 25/3 +h^2 = - 128/3

OpenStudy (zpupster):

then complete the square by taking one half coefficient of h and square it so 25/3*1/2 = 25/6 squaring gives 625/36 and add that to both sides giving 625/36 + 25/3 h +h^2 = -911/36

OpenStudy (zpupster):

this answer is negative and does not make sense so our answer is just because you can not have a negative volume so H=5

OpenStudy (zpupster):

or 5 8 16

OpenStudy (anonymous):

how did you get it?

OpenStudy (zpupster):

oh h=5 and l = 3(5) +1 and w= 3+5